Mr. Housenecht can do a piece of work in two and a half days and Mr. Warwick in three and one-third days. In how many days can both do it.
Answer by TimothyLamb(4379)

rate is work over time:
r = w/t
---
work can be anything such as 1 for one "job", or 1 for "a piece of work", or it can be a fraction or a multiple of a job, for example 0.75 of a job, or 3.83 jobs, etc.
---
each man's rate:
---
rate Housenecht:
---
two and a half days = 2+1/2 = 2.5 days
rateHousenecht = w/t
rateHousenecht = 1/(2+1/2)
rateHousenecht = 1/2.5
---
rate Warwick:
---
three and one-third days = 3+1/3 = 3.333333333
rateWarwick = w/t
rateWarwick = 1/(3+1/3)
rateWarwick = 1/3.333333333
---
the rates of the two men are additive:
r = rateHousenecht + rateWarwick
r = 1/2.5 + 1/3.333333333
r = 0.7
---
t = w/r
t = 1/0.7
t = 1.42857142857
---
answer:
together they can finish in about 1.43 days
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---